Solve the equation by completing the square.
a. Write irrational numbers in simplest radical form.
b. Describe each step as you go.
X2 + 4x – 8 = 0
Solve the equation by using the quadratic formula. Use √( ) and ± , if you want.
a. Write irrational numbers in simplest radical form.
b. Explain each step as you go.
X2 + 4x – 8 = 0





Answer by stanbon(75887)

Solve the equation by completing the square.
a. Write irrational numbers in simplest radical form.
b. Describe each step as you go.
X^2 + 4x – 8 = 0
x^2+4x+? = 8+?
x^2+4x+4 = 8+4
(x+2)^2 = 12
x+2 = 2sqrt(3) or x+2 = -2sqrt(3)
x = -2+2sqrt(3) or x = -2-2sqrt(3)
--------------------------------------------------
Solve the equation by using the quadratic formula. Use √( ) and ± , if you want.
a. Write irrational numbers in simplest radical form.
b. Explain each step as you go.
X2 + 4x – 8 = 0
---
x = [-4 +- sqrt(16-4*-8)]/2
x = [-4 +- sqrt(48)]/2
x = [-4 +- 4sqr(3)]/2
x = -2+2sqr(3) or x = -2-2sqrt(3)
======================================
Cheers,
Stan H.
==========